How do I find the height of this quadratic word problem?
If a ball is thrown straight up with an initial velocity of 48 feet per second, its height s after t seconds is given by the equation s=48t-16tsquared. Find the maximum height attained by the ball and the time it takes for the ball to return to earth.

Answer by khwang(438)

Apply complete square to the right of s(t), we have
Since s(t) = 48 t-16t^2 = -16(t^2 -3t + (3/2)^2) + 16(3/2)^2
= -16(t - 3/2)^2 + 36
We know -16(t - 3/2)^2 <= 0 for all t.
So, s(t) <= 36 and when t = 3/2 ,s(t) attains its maximum value 36.
Then to find how long it takes the ball returning the ground, we
have to solve s(t) = -16 t(t-3) = 0, we get t =3 or 0(inital).
[Or since we know that it takes 3/2 sec for the ball to reach its max height,
then it takes 2*3/2 =3 sec to return to earth.]
Hence, the maximum height of the ball is 36 ft and it takes 3 sec
for the ball to return to earth .